37: x^2 + 5x + 9 = 0
Answer by checkley75(3666)

X^2+5X=9=0
USING THE QUADRATIC EQUATION WE GET:
X=(-5+-SQRT[5^2-4*1*9])/281
X=(-5+-SQRT[25-36])/2
X=(-5+-SQRT-11)/2
X=(-5+-3.3177i)/2
X=-5/2+3.3177/2
X=-2.5+1.6583i ANSWER.
X=-5/2-3.3177i/2
X=-2.5-1.6583i ANSWER.